OpenStudy (blackops2luvr):

what is equal to the square root of 50

OpenStudy (blackops2luvr):

..

jimthompson5910 (jim_thompson5910):

Similar Example \[\large \sqrt{96} = \sqrt{16*6}\] \[\large \sqrt{96} = \sqrt{16}*\sqrt{6}\] \[\large \sqrt{96} = \sqrt{4^2}*\sqrt{6}\] \[\large \sqrt{96} = 4\sqrt{6}\] Notice how I factored 96 above into 16*6, and notice how 16 is a perfect square.

OpenStudy (blackops2luvr):

so its....

jimthompson5910 (jim_thompson5910):

Use the idea above to try to simplify \(\large \sqrt{50}\)

jimthompson5910 (jim_thompson5910):

what perfect squares are factors of 50?

OpenStudy (blackops2luvr):

dude im in 7th grade and in advanced math

jimthompson5910 (jim_thompson5910):

then you need to practice more, that's all

jimthompson5910 (jim_thompson5910):

if this is too advanced for you, then perhaps you need to either practice more or head to the class below this one